OpenStudy (anonymous):

Given the equation y − 3 = (x + 6) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (johnweldon1993):

Slope intercept form looks like \[y = mx + b\] you have y - 3 = x + 6 what can you do to get 'y' by itself on that left side of the equation?

OpenStudy (anonymous):

do you do multiplication?

OpenStudy (johnweldon1993):

not quite y - 3 = x + 6 you can add 3 to both sides of the equation... -3 + 3 = 0 so it cancels it out on that left side....so y - 3 = x + 6 +3 +3 y + 0 = x + 9 y = x + 9 do you get why we add instead of multiply?

OpenStudy (anonymous):

wow ok i get it now hahaha

OpenStudy (anonymous):

thank you so much

OpenStudy (johnweldon1993):

anytime :)